OpenStudy (anonymous):

Question (4/1Divided by 5/6) I got 24/5 but my calculator is getting 4 4/5 why is it 4 4/5

OpenStudy (anonymous):

\[\frac{4}{1}\div \frac{5}{6}=\frac{4}{1}\times \frac{6}{5}\]

OpenStudy (anonymous):

Calculators don't know the order of operations they are machine

OpenStudy (anonymous):

oh i see \[\frac{24}{5}=4\tfrac{4}{5}\]

OpenStudy (anonymous):

Is it a mixed one or is it 44/5 ?

OpenStudy (anonymous):

same answer one is an improper faction, the other a mixed number

OpenStudy (anonymous):

Why can't i input 24/5 instead of 4 4/5 if the black board is going to tell me im wrong..

OpenStudy (anonymous):

Since you guys can't tell me why it's 4 4/5. How to i change a improper fraction to a mixed number..